# First-principles Spin and Optical Properties of Vacancy Clusters in Lithium Fluoride

### v1

@@ Abstract
Figure 1: Photoluminescence spectrum of LiF, showing two broad emission peaks at 525nm and 650nm corresponding to the $F_{2}^{0}$-center and $F_{3}^{+}$-center respectively, with insets showing their structures.
@@ II.2 $F_{2}$-center
Our first principles calculations of the $F_{2}$-center and its charged counterparts, reveal two defect levels per spin-channel within the band gap (Figure 2b,c,d), along with several more non-degenerate unoccupied states above the CBM. This is consistent with electronic models of the $F_{2}$ center which predict a number of non-degenerate ground and excited states [51].
@@ II.3 $F_{3}$-center
Calculations of the defect states suggested a singlet ground state and excited state $e$ doublet for both $F_{3}^{0}$ [59] and $F_{3}^{+}$ [60] defects. Our computed electronic structures for the three $F_{3}$ charge states (Figure 2f,g,h) lack this double degeneracy, due to a breaking of the three-fold symmetry of the defect.
@@ III.1 Details of optical properties
Figure 5: Plots of $|\Psi|^{2}$ for the (a) highest occupied and (b) lowest unoccupied defect levels of the $F_{2}$-center, in the spin up (red) and spin down (blue) channels.
@@ III.3 Optical properties along the isoline
Table 1: HSE parameters $\alpha_{\text{HSE}}$ (exact exchange fraction) and $\mu_{\text{HSE}}$ (range separation parameter) satisfying generalized Koopman’s theorem for each defect, and the resulting defect level positions ($\Delta_{\text{HOMO}}$, $\Delta_{\text{LUMO}}$) relative to the conduction band. The spin states $\bm{S}$, absorption/emission energies ($E_{\text{a}}$, $E_{\text{e}}$)/wavelengths ($\lambda_{\text{a}}$, $\lambda_{\text{e}}$) compared with experimental values [38] ($\lambda^{\text{EX}}_{\text{a}}$, $\lambda^{\text{EX}}_{\text{e}}$), transition dipole moments (TDM), and emission radiative lifetimes ($\tau$, $\tau^{\text{EX}}$ [74]) are given for each electronic transition.

### v2

@@ Abstract
Figure 1: Photoluminescence spectrum of LiF, showing two broad emission peaks at $\sim$525 nm and $\sim$650 nm corresponding to the $F_{2}^{0}$-center and $F_{3}^{+}$-center respectively, with insets showing their structures.
@@ II.2 $F_{2}$-center
Our first principles calculations of the $F_{2}$-center and its charged counterparts, reveal two defect levels per spin-channel within the band gap (Figs. 2(b),(c),(d), along with several more non-degenerate unoccupied states above the CBM. This is consistent with electronic models of the $F_{2}$ center which predict a number of non-degenerate ground and excited states [51].
@@ II.3 $F_{3}$-center
Calculations of the defect states suggested a singlet ground state and excited state $e$ doublet for both $F_{3}^{0}$ [59] and $F_{3}^{+}$ [60] defects. Our computed electronic structures for the three $F_{3}$ charge states (Figs. 2(f),(g),(h) lack this double degeneracy, due to a breaking of the three-fold symmetry of the defect.
@@ III.1 Details of optical properties
Figure 5: Plots of $|\Psi|^{2}$ for the (a) higest occupied and (b) lowest unoccupied defect levels of the $F_{2}$-center, in the spin up (red) and spin down (blue) channels.
@@ III.3 Optical properties along the isoline
Table 1: HSE parameters $\alpha_{\text{HSE}}$ (exact exchange fraction) and $\mu_{\text{HSE}}$ (range separation parameter) satisfing generalized Koopman’s theorem for each defect, and the resulting defect level positions ($\Delta_{\text{HOMO}}$, $\Delta_{\text{LUMO}}$) relative to the conduction band. The spin states $\bm{S}$, absorption/emission energies ($E_{\text{a}}$, $E_{\text{e}}$)/wavelengths ($\lambda_{\text{a}}$, $\lambda_{\text{e}}$) compared with experimental values [38] ($\lambda^{\text{EX}}_{\text{a}}$, $\lambda^{\text{EX}}_{\text{e}}$), transition dipole moments (TDM), and emission radiative lifetimes ($\tau$, $\tau^{\text{EX}}$ [74]) are given for each electronic transition.